# CLEAR: Calibrated Learning for Epistemic and Aleatoric Risk

### v1

@@ E.3 Variant (c)
Table 19: Variant (c) NCIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or $<0.01$ are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive UACQR-P UACQR-S
ailerons 0.210 $\pm$ 0.020 0.231 $\pm$ 0.023 0.204 $\pm$ 0.021 0.239 $\pm$ 0.021 0.230 $\pm$ 0.021 0.188 $\pm$ 0.030 0.207 $\pm$ 0.017
airfoil 0.189 $\pm$ 0.013 0.255 $\pm$ 0.033 0.204 $\pm$ 0.016 0.199 $\pm$ 0.013 0.204 $\pm$ 0.014 0.317 $\pm$ 0.025 0.329 $\pm$ 0.033
allstate 0.278 $\pm$ 0.055 0.315 $\pm$ 0.066 0.280 $\pm$ 0.041 0.305 $\pm$ 0.070 0.319 $\pm$ 0.051 0.279 $\pm$ 0.047 0.296 $\pm$ 0.049
ca_housing 0.310 $\pm$ 0.009 0.418 $\pm$ 0.015 0.343 $\pm$ 0.011 0.329 $\pm$ 0.009 0.401 $\pm$ 0.018 0.368 $\pm$ 0.007 0.405 $\pm$ 0.007
computer 0.081 $\pm$ 0.004 0.104 $\pm$ 0.007 0.090 $\pm$ 0.004 0.084 $\pm$ 0.005 0.092 $\pm$ 0.005 0.087 $\pm$ 0.005 0.101 $\pm$ 0.001
concrete 0.259 $\pm$ 0.030 0.299 $\pm$ 0.053 0.263 $\pm$ 0.028 0.257 $\pm$ 0.031 0.265 $\pm$ 0.027 0.332 $\pm$ 0.056 0.344 $\pm$ 0.054
elevator 0.127 $\pm$ 0.007 0.193 $\pm$ 0.008 0.138 $\pm$ 0.009 0.151 $\pm$ 0.011 0.144 $\pm$ 0.010 0.165 $\pm$ 0.015 0.200 $\pm$ 0.010
energy_efficiency 0.042 $\pm$ 0.006 0.133 $\pm$ 0.051 0.049 $\pm$ 0.005 0.043 $\pm$ 0.005 0.049 $\pm$ 0.005 0.140 $\pm$ 0.014 0.138 $\pm$ 0.024
insurance 0.348 $\pm$ 0.079 0.421 $\pm$ 0.044 0.424 $\pm$ 0.074 0.343 $\pm$ 0.087 0.385 $\pm$ 0.112 0.291 $\pm$ 0.043 0.301 $\pm$ 0.034
kin8nm 0.322 $\pm$ 0.008 0.418 $\pm$ 0.017 0.327 $\pm$ 0.010 0.331 $\pm$ 0.010 0.341 $\pm$ 0.012 0.433 $\pm$ 0.009 0.445 $\pm$ 0.013
miami_housing 0.088 $\pm$ 0.002 0.114 $\pm$ 0.003 0.111 $\pm$ 0.007 0.094 $\pm$ 0.003 0.122 $\pm$ 0.007 0.102 $\pm$ 0.004 0.110 $\pm$ 0.004
naval_propulsion 1.6e-03 $\pm$ 4.9e-05 4.1e-03 $\pm$ 2.6e-04 1.6e-03 $\pm$ 4.8e-05 1.7e-03 $\pm$ 4.7e-05 1.8e-03 $\pm$ 5.8e-05 1.7e-03 $\pm$ 8.2e-05 3.6e-03 $\pm$ 8.5e-05
parkinsons 0.251 $\pm$ 0.012 0.366 $\pm$ 0.017 0.278 $\pm$ 0.013 0.259 $\pm$ 0.009 0.284 $\pm$ 0.013 0.273 $\pm$ 0.019 0.313 $\pm$ 0.018
powerplant 0.160 $\pm$ 0.007 0.176 $\pm$ 0.008 0.164 $\pm$ 0.008 0.171 $\pm$ 0.008 0.166 $\pm$ 0.007 0.191 $\pm$ 0.009 0.200 $\pm$ 0.009
qsar 0.349 $\pm$ 0.119 0.461 $\pm$ 0.155 0.358 $\pm$ 0.123 0.400 $\pm$ 0.139 0.393 $\pm$ 0.132 0.409 $\pm$ 0.134 0.449 $\pm$ 0.150
sulfur 0.109 $\pm$ 0.008 0.124 $\pm$ 0.008 0.108 $\pm$ 0.006 0.124 $\pm$ 0.008 0.115 $\pm$ 0.007 0.114 $\pm$ 0.010 0.126 $\pm$ 0.010
superconductor 0.192 $\pm$ 0.019 0.252 $\pm$ 0.026 0.206 $\pm$ 0.027 0.230 $\pm$ 0.023 0.297 $\pm$ 0.030 0.220 $\pm$ 0.023 0.232 $\pm$ 0.025
@@
Table 20: Variant (c) INTERVALSCORELOSS at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or $<0.01$ are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive UACQR-P UACQR-S
ailerons 7.7e-04 $\pm$ 2.4e-05 8.2e-04 $\pm$ 2.2e-05 7.9e-04 $\pm$ 2.9e-05 8.5e-04 $\pm$ 2.2e-05 9.4e-04 $\pm$ 3.0e-05 7.9e-04 $\pm$ 3.1e-05 7.9e-04 $\pm$ 2.3e-05
airfoil 8.567 $\pm$ 1.121 11.223 $\pm$ 0.608 10.339 $\pm$ 1.545 8.799 $\pm$ 1.034 10.361 $\pm$ 1.499 15.052 $\pm$ 1.025 15.999 $\pm$ 0.776
allstate 9.2e+03 $\pm$ 5.5e+02 9.6e+03 $\pm$ 4.1e+02 1.0e+04 $\pm$ 1.6e+03 1.1e+04 $\pm$ 7.8e+02 1.3e+04 $\pm$ 8.5e+02 9.0e+03 $\pm$ 5.9e+02 9.3e+03 $\pm$ 5.2e+02
ca_housing 2.2e+05 $\pm$ 7.3e+03 2.5e+05 $\pm$ 5.6e+03 2.5e+05 $\pm$ 8.4e+03 2.3e+05 $\pm$ 7.0e+03 2.9e+05 $\pm$ 9.7e+03 2.4e+05 $\pm$ 4.7e+03 2.4e+05 $\pm$ 4.6e+03
computer 10.973 $\pm$ 0.906 12.139 $\pm$ 1.009 12.826 $\pm$ 0.947 11.204 $\pm$ 0.996 13.130 $\pm$ 1.051 11.757 $\pm$ 0.463 12.103 $\pm$ 0.323
concrete 26.775 $\pm$ 2.471 29.923 $\pm$ 3.423 30.402 $\pm$ 4.655 26.855 $\pm$ 2.535 30.720 $\pm$ 4.505 33.077 $\pm$ 2.941 33.990 $\pm$ 1.797
elevator 0.010 $\pm$ 0.000 0.014 $\pm$ 0.000 0.012 $\pm$ 0.000 0.011 $\pm$ 0.000 0.013 $\pm$ 0.000 0.014 $\pm$ 0.000 0.015 $\pm$ 0.000
energy_efficiency 2.338 $\pm$ 0.553 5.406 $\pm$ 0.356 2.760 $\pm$ 0.823 2.397 $\pm$ 0.573 2.756 $\pm$ 0.822 6.214 $\pm$ 0.408 6.354 $\pm$ 0.480
insurance 3.4e+04 $\pm$ 6.5e+03 3.1e+04 $\pm$ 2.7e+03 3.2e+04 $\pm$ 4.2e+03 3.5e+04 $\pm$ 5.2e+03 3.3e+04 $\pm$ 3.3e+03 2.7e+04 $\pm$ 3.7e+03 2.7e+04 $\pm$ 2.8e+03
kin8nm 0.534 $\pm$ 0.014 0.673 $\pm$ 0.021 0.544 $\pm$ 0.015 0.552 $\pm$ 0.017 0.591 $\pm$ 0.022 0.705 $\pm$ 0.013 0.723 $\pm$ 0.013
miami_housing 2.9e+05 $\pm$ 1.1e+04 3.3e+05 $\pm$ 8.9e+03 5.9e+05 $\pm$ 3.4e+04 3.0e+05 $\pm$ 9.6e+03 6.1e+05 $\pm$ 3.4e+04 3.5e+05 $\pm$ 1.0e+04 3.6e+05 $\pm$ 1.0e+04
naval_propulsion 3.2e-03 $\pm$ 8.6e-05 7.6e-03 $\pm$ 1.6e-04 3.9e-03 $\pm$ 1.2e-04 3.4e-03 $\pm$ 9.6e-05 4.6e-03 $\pm$ 9.6e-05 7.5e-03 $\pm$ 7.5e-04 6.6e-03 $\pm$ 1.3e-04
parkinsons 15.156 $\pm$ 0.516 19.920 $\pm$ 0.599 17.322 $\pm$ 0.628 15.312 $\pm$ 0.459 17.596 $\pm$ 0.621 14.384 $\pm$ 0.700 15.978 $\pm$ 0.779
powerplant 16.242 $\pm$ 0.989 17.038 $\pm$ 1.138 16.671 $\pm$ 0.987 16.846 $\pm$ 1.029 16.874 $\pm$ 0.975 18.078 $\pm$ 0.906 18.725 $\pm$ 0.822
qsar 3.850 $\pm$ 0.177 4.528 $\pm$ 0.221 4.083 $\pm$ 0.194 4.278 $\pm$ 0.129 4.392 $\pm$ 0.203 4.102 $\pm$ 0.230 4.333 $\pm$ 0.182
sulfur 0.143 $\pm$ 0.005 0.156 $\pm$ 0.010 0.213 $\pm$ 0.013 0.146 $\pm$ 0.005 0.228 $\pm$ 0.013 0.153 $\pm$ 0.010 0.158 $\pm$ 0.012
superconductor 39.048 $\pm$ 1.591 44.219 $\pm$ 1.421 45.523 $\pm$ 3.375 45.127 $\pm$ 1.849 63.336 $\pm$ 2.769 40.545 $\pm$ 1.386 42.122 $\pm$ 1.257
@@
ailerons 1.05 [0.14:3.46] 0.98 [0.89:1.58]
airfoil 0.76 [0.16:5.93] 1.72 [0.22:4.56]
allstate 0.23 [0.00:15.89] 1.11 [0.15:1.94]
ca_housing 1.83 [1.01:3.70] 0.72 [0.42:0.98]
computer 2.29 [1.61:8.91] 0.73 [0.21:0.95]
concrete 1.95 [0.06:100.00] 0.84 [0.01:16.28]
elevator 0.95 [0.64:1.39] 1.03 [0.84:1.24]
energy_efficiency 0.12 [0.01:100.00] 7.27 [0.02:22.89]
insurance 5.06 [0.14:100.00] 0.38 [0.02:28.07]
kin8nm 1.98 [0.81:4.96] 0.75 [0.50:0.94]
miami_housing 6.88 [3.92:20.81] 0.24 [0.09:0.40]
naval_propulsion 16.32 [12.50:21.65] 0.64 [0.52:0.76]
parkinsons 1.18 [0.68:8.82] 1.26 [0.21:1.91]
powerplant 1.08 [0.82:3.62] 1.07 [0.47:1.27]
qsar 0.73 [0.41:1.32] 1.36 [1.04:2.13]
sulfur 2.49 [1.53:8.08] 0.69 [0.26:0.91]
superconductor 0.84 [0.37:1.79] 1.09 [0.69:1.30]
E.4 Comparing all variants of CLEAR

### v2

@@ F.3 Variant (c)
Table 25: Variant (c) NCIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
ailerons 0.210 $\pm$ 0.020 0.231 $\pm$ 0.023 0.204 $\pm$ 0.021 0.239 $\pm$ 0.021 0.230 $\pm$ 0.021 0.209 $\pm$ 0.020 0.210 $\pm$ 0.019
airfoil 0.188 $\pm$ 0.014 0.258 $\pm$ 0.034 0.204 $\pm$ 0.017 0.199 $\pm$ 0.013 0.204 $\pm$ 0.014 0.190 $\pm$ 0.012 0.188 $\pm$ 0.010
allstate 0.277 $\pm$ 0.056 0.316 $\pm$ 0.065 0.281 $\pm$ 0.041 0.305 $\pm$ 0.070 0.319 $\pm$ 0.051 0.271 $\pm$ 0.054 0.270 $\pm$ 0.053
ca_housing 0.311 $\pm$ 0.010 0.419 $\pm$ 0.014 0.343 $\pm$ 0.010 0.329 $\pm$ 0.009 0.401 $\pm$ 0.018 0.309 $\pm$ 0.009 0.309 $\pm$ 0.009
computer 0.081 $\pm$ 0.004 0.104 $\pm$ 0.007 0.090 $\pm$ 0.004 0.084 $\pm$ 0.005 0.092 $\pm$ 0.005 0.079 $\pm$ 0.004 0.080 $\pm$ 0.004
concrete 0.259 $\pm$ 0.032 0.303 $\pm$ 0.054 0.263 $\pm$ 0.029 0.257 $\pm$ 0.031 0.265 $\pm$ 0.027 0.258 $\pm$ 0.033 0.259 $\pm$ 0.032
elevator 0.126 $\pm$ 0.008 0.192 $\pm$ 0.008 0.138 $\pm$ 0.008 0.151 $\pm$ 0.011 0.144 $\pm$ 0.010 0.127 $\pm$ 0.007 0.127 $\pm$ 0.007
energy_efficiency 0.041 $\pm$ 0.006 0.116 $\pm$ 0.015 0.048 $\pm$ 0.005 0.043 $\pm$ 0.005 0.049 $\pm$ 0.005 0.042 $\pm$ 0.006 0.042 $\pm$ 0.006
insurance 0.345 $\pm$ 0.078 0.413 $\pm$ 0.043 0.424 $\pm$ 0.074 0.343 $\pm$ 0.087 0.385 $\pm$ 0.112 0.367 $\pm$ 0.100 0.407 $\pm$ 0.137
kin8nm 0.323 $\pm$ 0.008 0.420 $\pm$ 0.018 0.327 $\pm$ 0.008 0.331 $\pm$ 0.010 0.341 $\pm$ 0.012 0.325 $\pm$ 0.009 0.324 $\pm$ 0.009
miami_housing 0.088 $\pm$ 0.002 0.113 $\pm$ 0.004 0.111 $\pm$ 0.007 0.094 $\pm$ 0.003 0.122 $\pm$ 0.007 0.077 $\pm$ 0.003 0.076 $\pm$ 0.003
naval_propulsion 1.6e-03 $\pm$ 5.0e-05 4.0e-03 $\pm$ 2.4e-04 1.6e-03 $\pm$ 4.1e-05 1.7e-03 $\pm$ 4.7e-05 1.8e-03 $\pm$ 5.8e-05 1.6e-03 $\pm$ 5.2e-05 1.6e-03 $\pm$ 6.2e-05
parkinsons 0.250 $\pm$ 0.012 0.366 $\pm$ 0.017 0.279 $\pm$ 0.013 0.259 $\pm$ 0.009 0.284 $\pm$ 0.013 0.251 $\pm$ 0.011 0.250 $\pm$ 0.011
powerplant 0.161 $\pm$ 0.007 0.176 $\pm$ 0.009 0.163 $\pm$ 0.008 0.171 $\pm$ 0.008 0.166 $\pm$ 0.007 0.161 $\pm$ 0.007 0.161 $\pm$ 0.007
qsar 0.348 $\pm$ 0.118 0.465 $\pm$ 0.156 0.361 $\pm$ 0.124 0.407 $\pm$ 0.142 0.395 $\pm$ 0.134 0.359 $\pm$ 0.124 0.354 $\pm$ 0.122
sulfur 0.109 $\pm$ 0.008 0.124 $\pm$ 0.008 0.108 $\pm$ 0.006 0.124 $\pm$ 0.008 0.115 $\pm$ 0.007 0.104 $\pm$ 0.007 0.102 $\pm$ 0.007
superconductor 0.192 $\pm$ 0.019 0.251 $\pm$ 0.025 0.206 $\pm$ 0.027 0.232 $\pm$ 0.025 0.297 $\pm$ 0.029 0.191 $\pm$ 0.019 0.191 $\pm$ 0.019
@@
Table 26: Variant (c) Interval Score Loss at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
ailerons 7.7e-04 $\pm$ 2.3e-05 8.2e-04 $\pm$ 2.2e-05 7.9e-04 $\pm$ 2.9e-05 8.5e-04 $\pm$ 2.2e-05 9.4e-04 $\pm$ 3.0e-05 7.7e-04 $\pm$ 2.3e-05 7.7e-04 $\pm$ 2.3e-05
airfoil 8.569 $\pm$ 1.099 11.217 $\pm$ 0.549 10.341 $\pm$ 1.548 8.799 $\pm$ 1.034 10.361 $\pm$ 1.499 8.675 $\pm$ 1.116 8.624 $\pm$ 1.105
allstate 9.2e+03 $\pm$ 5.5e+02 9.6e+03 $\pm$ 4.0e+02 1.0e+04 $\pm$ 1.6e+03 1.1e+04 $\pm$ 7.8e+02 1.3e+04 $\pm$ 8.5e+02 9.2e+03 $\pm$ 5.9e+02 9.3e+03 $\pm$ 5.9e+02
ca_housing 2.2e+05 $\pm$ 7.4e+03 2.5e+05 $\pm$ 5.7e+03 2.5e+05 $\pm$ 8.1e+03 2.3e+05 $\pm$ 7.0e+03 2.9e+05 $\pm$ 9.7e+03 2.2e+05 $\pm$ 7.1e+03 2.2e+05 $\pm$ 7.0e+03
computer 10.978 $\pm$ 0.911 12.134 $\pm$ 1.001 12.827 $\pm$ 0.945 11.204 $\pm$ 0.996 13.130 $\pm$ 1.051 10.986 $\pm$ 0.873 11.030 $\pm$ 0.866
concrete 26.928 $\pm$ 2.529 29.955 $\pm$ 3.544 30.449 $\pm$ 4.693 26.855 $\pm$ 2.535 30.720 $\pm$ 4.505 26.882 $\pm$ 2.568 26.962 $\pm$ 2.581
elevator 0.010 $\pm$ 0.000 0.014 $\pm$ 0.000 0.012 $\pm$ 0.000 0.011 $\pm$ 0.000 0.013 $\pm$ 0.000 0.010 $\pm$ 0.000 0.010 $\pm$ 0.000
energy_efficiency 2.348 $\pm$ 0.554 5.407 $\pm$ 0.418 2.759 $\pm$ 0.822 2.397 $\pm$ 0.573 2.756 $\pm$ 0.822 2.392 $\pm$ 0.564 2.395 $\pm$ 0.588
insurance 3.4e+04 $\pm$ 6.4e+03 3.1e+04 $\pm$ 2.7e+03 3.2e+04 $\pm$ 4.3e+03 3.5e+04 $\pm$ 5.2e+03 3.3e+04 $\pm$ 3.3e+03 3.4e+04 $\pm$ 6.2e+03 3.6e+04 $\pm$ 5.4e+03
kin8nm 0.534 $\pm$ 0.014 0.673 $\pm$ 0.021 0.544 $\pm$ 0.015 0.552 $\pm$ 0.017 0.591 $\pm$ 0.022 0.536 $\pm$ 0.013 0.535 $\pm$ 0.013
miami_housing 2.9e+05 $\pm$ 1.1e+04 3.3e+05 $\pm$ 9.5e+03 5.9e+05 $\pm$ 3.4e+04 3.0e+05 $\pm$ 9.6e+03 6.1e+05 $\pm$ 3.4e+04 3.2e+05 $\pm$ 1.9e+04 3.2e+05 $\pm$ 1.9e+04
naval_propulsion 3.2e-03 $\pm$ 8.4e-05 7.6e-03 $\pm$ 1.7e-04 3.9e-03 $\pm$ 1.3e-04 3.4e-03 $\pm$ 9.6e-05 4.6e-03 $\pm$ 9.6e-05 3.2e-03 $\pm$ 9.3e-05 3.6e-03 $\pm$ 1.3e-04
parkinsons 15.149 $\pm$ 0.518 19.899 $\pm$ 0.576 17.316 $\pm$ 0.634 15.312 $\pm$ 0.459 17.596 $\pm$ 0.621 15.150 $\pm$ 0.499 15.161 $\pm$ 0.537
powerplant 16.243 $\pm$ 0.985 17.062 $\pm$ 1.128 16.673 $\pm$ 0.987 16.846 $\pm$ 1.029 16.874 $\pm$ 0.975 16.231 $\pm$ 0.984 16.231 $\pm$ 0.983
qsar 3.877 $\pm$ 0.176 4.521 $\pm$ 0.225 4.086 $\pm$ 0.193 4.334 $\pm$ 0.125 4.390 $\pm$ 0.205 3.927 $\pm$ 0.146 3.892 $\pm$ 0.151
sulfur 0.143 $\pm$ 0.005 0.156 $\pm$ 0.010 0.213 $\pm$ 0.013 0.146 $\pm$ 0.005 0.228 $\pm$ 0.013 0.143 $\pm$ 0.005 0.145 $\pm$ 0.006
superconductor 39.104 $\pm$ 1.604 44.230 $\pm$ 1.428 45.479 $\pm$ 3.358 45.456 $\pm$ 2.039 63.347 $\pm$ 2.824 39.224 $\pm$ 1.613 39.119 $\pm$ 1.642
@@
ailerons 0.89 [0.17:4.47] 0.98 [0.87:1.58]
airfoil 0.76 [0.23:5.25] 1.70 [0.25:4.30]
allstate 0.21 [0.00:14.06] 1.14 [0.16:1.93]
ca_housing 1.90 [1.14:5.08] 0.70 [0.33:0.93]
computer 2.21 [1.63:8.87] 0.74 [0.21:0.96]
concrete 1.33 [0.06:100.00] 1.65 [0.01:16.73]
elevator 0.94 [0.67:1.43] 1.04 [0.82:1.22]
energy_efficiency 0.12 [0.01:100.00] 7.96 [0.02:24.25]
insurance 5.07 [0.14:100.00] 0.38 [0.02:27.45]
kin8nm 1.87 [1.64:3.82] 0.76 [0.58:0.83]
miami_housing 7.06 [4.84:19.45] 0.23 [0.09:0.33]
naval_propulsion 16.16 [12.48:21.17] 0.63 [0.53:0.76]
parkinsons 1.16 [0.72:8.42] 1.26 [0.22:1.84]
powerplant 1.05 [0.73:3.26] 1.05 [0.51:1.36]
qsar 0.51 [0.32:1.70] 1.45 [0.96:2.37]
sulfur 2.41 [1.53:7.16] 0.71 [0.29:0.91]
superconductor 0.91 [0.34:1.39] 1.07 [0.80:1.34]
F.4 Comparing Variants of CLEAR